Determine whether the vectors a=2i-3j+5k,b=3i+j-2k and c=i-7j+12k are linearly dependent or linearly independent
1
Expert's answer
2019-10-04T10:56:42-0400

"det\begin{bmatrix}
 2 & -3&5 \\
 3 & 1&-2\\
 1&-7&12
\end{bmatrix}=2*1*12+3*5*(-7)+(-3)*(-2)*1-\\-
1*1*5-2*(-7)*(-2)-3*(-3)*12=0."

So, the vectors are linearly dependent.